OpenStudy (anonymous):

what is the value of What is the value of f[g(6)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

calculate first g(6) then f of this value

OpenStudy (anonymous):

Since, g(x) = 2x - 5 i.e. g(6) = 2*6 - 5=12-5=7 i.e. g(6) =7 since f(x) = 2x + 1 Now substituting x = g(6) in f(x) = 2x + 1 we find f(g(6)) = 2g(6) + 1 i.e. f(g(6)) = 2*7 + 1=14+1=15 Thus f[g(6)] = 15